As a financial analyst at Glencolin International (GI) you have been asked to evaluate two capital investment alternatives submitted by the production department of the firm. Before beginning your analysis, you note that company policy has set the cost of capital at 15 percent for all proposed projects. As a small business, GI pays corporate taxes at the rate of 35 percent. The proposed capital project calls for developing new computer software to facilitate partial automation of production in GI's plant. Alternative A has initial software development costs projected at $194,000, while Alternative B would cost $336,000. Software development costs would be capitalized and qualify for a capital cost allowance (CCA) rate of 30 percent. In addition, IT would hire a software consultant under either alternative to assist in making the decision whether to invest in the project for a fee of $17,000 and this cost would be expensed when it is incurred. To recover its costs, GI's IT department would charge the production department for the use of computer time at the rate of $394 per hour and estimates that it would take 182 hours of computer time per year to run the new software under either alternative. GI owns all its computers and does not currently operate them at capacity. The information technology (IT) plan calls for this excess capacity to continue in the future. For security reasons, it is company policy not to rent excess computing capacity to outside users. If the new partial automation of production is put in place, expected savings in production cost (before tax) are projected as follows:


Year--- Alternative A ------- Alternative B
1 -----------$86,000 ----------------$118,000
2 -----------$86,000 ----------------$130,000
3 -----------$67,000---------------- $106,000
4 -----------$56,000---------------- $98,000
5 -----------$39,000---------------- $59,000

As a financial analyst at Glencolin International (GI) you have been asked to revisit your analysis of the two capital investment alternatives submitted by the production department of the firm. (Detailed discussion of these alternatives is in the Mini Case at the end of Chapter 10.) The CFO is concerned that the analysis to date has not really addressed the risk in this project. Your task is to employ scenario and sensitivity analysis to explore how your original recommendation might change when subjected to a number of “what-ifs.” In your discussions with the CFO, the CIO and the head of the production department, you have pinpointed two key inputs to the capital budgeting decision: initial software development costs and expected savings in production costs (before tax). By properly designing the contract for software development, you are confident that initial software costs for each alternative can be kept in a range of plus or minus 15 percent of the original estimates. Savings in production costs are less certain because the software will involve new technology that has not been implemented before. An appropriate range for these costs is plus or minus 40 percent of the original estimates. As the capital budgeting analyst, you are required to answer the following in your memo to the CFO:

a. a) Conduct sensitivity analysis to determine which of the two inputs has a greater input on the choice between the two projects.

b. b) Conduct scenario analysis to assess the risks of each alternative in turn. What are your conclusions?

c. c) Explain what your sensitivity and scenario analyses tell you about your original recommendations.
Answer #1
Software Consultant fees of $17,000 is sunkcost and not relevant for this analysis
Present Value (PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=Discount Rate=cost of capital =15% 0.15
N=Year of Cash Flow
ALTERNATIVE A
year 0 1 2 3 4 5
A Initial Cash Flow ($194,000)
B Before tax Savings $86,000 $86,000 $67,000 $56,000 $39,000
C=B*(1-0.35) After tax savings $55,900 $55,900 $43,550 $36,400 $25,350
D AnnualDepreciation $58,200 $40,740 $28,518 $19,963 $13,974
E= Accumulated depreciation $58,200 $98,940 $127,458 $147,421 $161,394
F=194000-E Year End Book value $135,800 $95,060 $66,542 $46,579 $32,606
G=D*0.35 Depreciation Tax Shield $20,370 $14,259 $9,981 $6,987 $4,891
H=A+C+G Net Cash Flow ($194,000) $76,270 $70,159 $53,531 $43,387 $30,241 SUM
I=H/(1.15^N) Present value of Cash Flow ($194,000) $66,322 $53,050 $35,198 $24,807 $15,035 $411
NET PRESENT VALUE OF ALTERNATIVE -A $411
ALTERNATIVE B
year 0 1 2 3 4 5
A Initial Cash Flow ($336,000)
B Before tax Savings $118,000 $130,000 $106,000 $98,000 $59,000
C=B*(1-0.35) After tax savings $76,700 $84,500 $68,900 $63,700 $38,350
D AnnualDepreciation $100,800 $70,560 $49,392 $34,574 $24,202
E= Accumulated depreciation $100,800 $171,360 $220,752 $255,326 $279,528
F=336000-E Year End Book value $235,200 $164,640 $115,248 $80,674 $56,472
G=D*0.35 Depreciation Tax Shield $35,280 $24,696 $17,287 $12,101 $8,471
H=A+C+G Net Cash Flow ($336,000) $111,980 $109,196 $86,187 $75,801 $46,821 SUM
I=H/(1.15^N) Present value of Cash Flow ($336,000) $97,374 $82,568 $56,669 $43,339 $23,278 ($32,771)
NET PRESENT VALUE OF ALTERNATIVE -B ($32,771)
SENSITIVITY ANALYSIS:
10% DECREASE IN AnnualSavings
ALTERNATIVE A
year 0 1 2 3 4 5
A Initial Cash Flow ($194,000)
B Before tax Savings $77,400 $77,400 $60,300 $50,400 $35,100
C=B*(1-0.35) After tax savings $50,310 $50,310 $39,195 $32,760 $22,815
D AnnualDepreciation $58,200 $40,740 $28,518 $19,963 $13,974
E= Accumulated depreciation $58,200 $98,940 $127,458 $147,421 $161,394
F=194000-E Year End Book value $135,800 $95,060 $66,542 $46,579 $32,606
G=D*0.35 Depreciation Tax Shield $20,370 $14,259 $9,981 $6,987 $4,891
H=A+C+G Net Cash Flow ($194,000) $70,680 $64,569 $49,176 $39,747 $27,706 SUM
I=H/(1.15^N) Present value of Cash Flow ($194,000) $61,461 $48,823 $32,334 $22,725 $13,775 ($14,881)
NET PRESENT VALUE OF ALTERNATIVE -A ($14,881)
ALTERNATIVE B
year 0 1 2 3 4 5
A Initial Cash Flow ($336,000)
B Before tax Savings $106,200 $117,000 $95,400 $88,200 $53,100
C=B*(1-0.35) After tax savings $69,030 $76,050 $62,010 $57,330 $34,515
D Annual SCENARIO ANALYSIS: BEST CASE SCENARIO Initial cost reduced by 15%,Annual savings increased by 40% ALTERNATIVE A ear Initial Cash Flow Before tax Savings ($164,900) $120,400$120,400 $93,800 $78,400 $54,600 $78,260 $78,260 $60,970 $50,960 $35,490 $49,470 $34,629 $24,240 $16,968 $11,878 $49,470 $84,099 $108,339 $125,308 $137,185 $115,430$80,801 $56,561 $39,592 $27,715 $17,315 $12,120 $8,484 $5,939 $4,157 C B*(1-0.35) After tax savings AnnualDepreciation Accumulated depreciation F-164900-E Year End Book value G=D*0.35 Depreciation Tax Shield HEA+C+G Net Cash Flow I=H/(1.15aN) Present value of Cash Flow ($164,900) $95,575 $90,380 $69,454 $56,899 $39,647 SUM ($164,900) $83,108 $68,340 $45,667$32,532 $19,712 $84,460 NET PRESENT VALUE OF ALTERNATIVE-A $84,460BEST CASE SCENARIO ALTERNATIVE B ear Initial Cash Flow Before tax Savings ($285,600) $165,200$182,000 $148,400 $137,200 $82,600 $107,380$118,300 $96,460 $89,180 $53,690 $85,680 $59,976 $41,983 $29,388 $20,572 $85,680 $145,656 $187,639 $217,027 $237,599 $199,920$139,944 $97,961 $68,573 $48,001 $29,988 $20,992 $14,694 $10,286$7,200 C B*(1-0.35) After tax savings AnnualDepreciation Accumulated depreciation F-285600-E Year End Book value G D*0.35 Depreciation Tax Shield H-A+C+G Net Cash Flow I=H/(1.15^N) Present value of Cash Flow ($285,600) $137,368 $139,292 $111,154 $99,466 $60,890 SUM ($285,600) $119,450 $105,324 $73,086 $56,870 $30,273 $99,404 NET PRESENT VALUE OF ALTERNATIVE -B $99,404ALTERNATIVE A year Initial Cash Flow Before tax Savings ($223,100) $51,600 $51,600 $40,200 $33,600 $23,400 33,540 $33,540 $26,130 $21,840 $15,210 $66,930 $46,851 $32,796 $22,957 $16,070 $66,930 $113,781 $146,577 $169,534 $185,604 156,170 $109,319 $76,523 $53,566 $37,496 23,426 $16,398 $11,478 $8,035 $5,624 C-B (1-0.35) After tax savings AnnualDepreciation Accumulated depreciation F-223100-E Year End Book value G=D*0.35 Depreciation Tax Shield HEA+C+G Net Cash Flow I-H/(1.15AN) Present value of Cash Flow ($223,100) $56,966 $49,938 $37,608 $29,875 $20,834 SUM ($223,100) $49,535 37,760 $24,728 $17,081 $10,358 (83,637) NET PRESENT VALUE OF ALTERNATIVE -A ($83,637)WORST CASE SCENARIO ALTERNATIVE B ear Initial Cash Flow Before tax Savings 386,400 $70,800 $78,000 $63,600 $58,800 $35,400 $46,020 $50,700 $41,340 $38,220 $23,010 115,920 $81,144 $56,801 $39,761 $27,832 $115,920 $197,064 $253,865 $293,625 $321,458 $270,480$189,336 $132,535 $92,775 $64,942 40,572 $28,400 $19,880 $13,916 $9,741 C-B*(1-0.35) Aftertax savings AnnualDepreciation Accumulated depreciation F-386400-E Year End Book value G=D*0.35 Depreciation Tax Shield HEA+C+G Net Cash Flow ($386,400) $86,592 $79,100 $61,220 $52,136 $32,751 SUM H/1.15 N) Present value of Cash Flow ($386,400) $75,297 $59,811 $40,253 $29,809 $16,283 ($164,946) NET PRESENTVALUE OF ALTERNATIVE -B ($164,946)Depreciation $100,800 $70,560 $49,392 $34,574 $24,202
E= Accumulated depreciation $100,800 $171,360 $220,752 $255,326 $279,528
F=336000-E Year End Book value $235,200 $164,640 $115,248 $80,674 $56,472
G=D*0.35 Depreciation Tax Shield $35,280 $24,696 $17,287 $12,101 $8,471
H=A+C+G Net Cash Flow ($336,000) $104,310 $100,746 $79,297 $69,431 $42,986 SUM
I=H/(1.15^N) Present value of Cash Flow ($336,000) $90,704 $76,178 $52,139 $39,697 $21,372 ($55,909)
NET PRESENT VALUE OF ALTERNATIVE -B ($55,909)